Evening shift August 20, 99

B. Cox (cox@fnal.gov)
Sat, 21 Aug 1999 00:12:35 -0500 (CDT)

Shift Summary for Evening Shift, Aug. 20,/99

On Shift: Brad, Hogan

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E832 data (special runs not included): 6.25
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

15:30 - 15:52 No beam

22:45 - 23: Memory bank DDD1 not responding
========================================================================
Runs ended this shift:

Run 14287
Run 14288
Run 14289
Run 14290
Run 14291

=======================================================================

Tapes taken to FCC this shift

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

15:30 Jason H. replaces a AZ latch in Crate 53 to
try to solve the begin of run ktevana errors.
Did not improve the situation. Will live with
occasional ktevana errors.

15:52 Beam returns. Contiue run 14287.

22:45 DDD1 veto comes on. Investigate.

22:50 Decide to end run 14287

22:50 - 23:15 Reset all DC2's in plane 1
Reset the VME167

Remove tapes QK*294-296, QKB103
Mount tapes QK*297-299, QKB104

Initialize for run 14288

23:15 Start Run 14288
Continue to have hangs on Stage3 initialization

23:15-23:50 Try various things. Start Run 14289. Still hangs on
stage 3 initialization

Check heartbeat and isolate problems
to DC2 of stream 3. Cycle power supply on back side of
VME crate of plane 1.

23:55 Start Run 14290. Damp summary process did not connect
to the dis server.

Try again. Start Run 14291.

========================================================================
Description of Shift:

Smooth running up to DDD1 problems

========================================================================
Follow up Notes on Previous Shift